Circular No. 6725

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997de IN NGC 6769 Alexander Wassilieff, Palmerston North, New Zealand, reports his dis-

covery, on two CCD images taken with a 0.36-m f/4 reflector (+ SBIG ST-7 camera) on Aug. 27, of an apparent supernova (mag ∼ 18.2) located 14′′.5 west and 20′′.6 north of the center of NGC 6769 (α = 19^{h}18^{m}.4, δ = −60^{o}31′, equinox 2000.0). The star does not appear on a similar image taken on 1994 July 9, nor does it appear on the Digital Sky Survey. B. Schmidt, Mount Stromlo and Siding Spring Observatories, reports via R. Evans that Brendan Downs (at Camp Duckadang, Queensland) independently found SN 1997de with a 0.31-m telescope (+ CCD) on Aug. 27.7, the offset noted as 16′′ west and 22′′ north of the center of NGC 6769.

N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports that spectrophotometry (range 320–900 nm; resolution 900) was obtained by J. Maza on Aug. 28.0 UT with the CTIO 4-m telescope. M. Phillips and Suntzeff find the spectrum of SN 1997de to be that of a type-Ia event, ∼ 4 weeks after maximum. There is significant Na I absorption (0.23 nm) at 3900 km/s (at the redshift of the host galaxy).

SUPERNOVA 1997dd IN NGC 6060 Suntzeff also reports that spectrophotometry (obtained as above) of SN

1997dd leads Phillips and himself to deduce this object to be a peculiar type-II supernova. Hα is clearly present but weak, and the redward emission is particularly weak. Hβ is extremely weak. Strong Ca II absorption at 390 and 840 nm is present, as is the absorption feature (Na I/He) at 580 nm. The spectrum appears to be intermediate between types II-P and Ib, and it is very similar to the spectrum SN 1987K, 7–9 days after maximum. SN 1997dd is thus expected to evolve rapidly into a type-Ib/c supernova.

GRB 970815 M. T. Adams, M. Ward, F. Ma, A. Howell, L. Wang, and J. C. Wheeler,

McDonald Observatory, report: “The field of GRB 970815 was observed with the primary-focus camera on the McDonald Observatory 0.76-m telescope. The images cover a field 46′ × 46′, centered on the RXTE burst position. The limiting magnitudes of the images are: Aug. 17.2 UT, R = 20.2, I = 19.9; 21.2, R = 20.4, I = 20.5; 25.2, I = 21; and 27.2, V = 21, I = 21. No objects brighter than R = 19.0 and I = 19.5 showed variability > 0.4 mag in the entire field during these observations.”

1997 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green